Givens T (N) m (kg) Viy (m/s) ry (m) 78 20 Scenario 1 Scenario 2 Trn A. What is the weight of the object? S1. S2. B. What is ay? S1. S2 C. what is rfv when t = 4 s? S1. S2 woll D. What is vjy when t 4 s? S1 S2.
Answer #1

[A] In all the questions I'll take g=10

Weight is same in both situations.

W=mg=80 extup{ N}

[B]

S1

T -mgmdy

ay = T/m-g =-0.25 m/s

which means the block is moving down.

S2

I'll assume downward this time.

mg-T cos 45° = ma

а,-у-1-cos 45° :: 3.1 m/s* ?7m

[C]

We already calculated the net downwards acceleration

S1

Use the following, for initial velocity 0,

r_{fy}=r_{iy}+rac{1}{2}a_yt^2

20--× 0.25 × 42-18 m rf,-

S2

Similarly, here

Yv = 20-2 x 3.1 × 4 2 4.85 m fy

[D]

For velocity

v_{fy}=v_{iy}+a_yt

S1

vfy0 0.25 x 4-1 m/:s

S2

12.4 m/s